Two cars are leaving Phils house at precisely 6:20 PM. The first car, driven by Phil is headed north. The second car, driven by Roman, is headed south. Phil, because he is not a good citizen, is going 20 MPH faster than Roman. If, at 7:50, they are exactly 168 miles apart, how fast is each going?
Answer by checkley75(3666)

BOTH CARS TRAVEL FOR 1.5 HOURS. THERE SPEEDS ARE X & (X+20). THEIR TOTAL DISTANCE IS 168 MILES THUS WE HAVE THE FOLLOWING FORMULA:
1.5X+1.5(X+20)=168
1.5X+1.5X+30=168
3X=168-30
3X=138
X=138/3
X=46 MPH FOR THE SLOWER DRIVER.
46+20=66 MPH FOR THE FASTER DRIVER.
PROOF
1.5*46+1.5*66=168
69+99=168
168=168